A coffee maker is rated at 1260 W, a toaster at 1060 W, and a waffle maker at 1460 W. The three appliances are connected in parallel to a common 120 V household circuit.

(a) What is the current in each appliance when operating independently?

coffee maker A
toaster A
waffle maker A


(b) What total current is delivered to the appliances when all are operating simultaneously?
A

(c) Is a 15-A circuit breaker sufficient in this situation? Explain.
Answer #1

a)
for coffee maker,

Apply, Power, P = V*I

==> I = P/V

= 1260/120

= 10.5 A

for toaster,

I = P/V

= 1060/120

= 8.83 A

for waffle maker,

I = P/V

= 1460/120

= 12.17 A


b) Itotal = 10.5 +8.83 + 12.17

= 31.5 A

c) No. The circuit needs 31.5 A circuit breaker